# Aligning to Thousands of Preferences via System Message Generalization

### v1

@@ Abstract
Although humans inherently have diverse values, current large language model (LLM) alignment methods often assume that aligning LLMs with the general public’s preferences is optimal. A major challenge in adopting a more individualized approach to LLM alignment is its lack of scalability, as it involves repeatedly acquiring preference data and training new reward models and LLMs for each individual’s preferences. To address these challenges, we propose a new paradigm where users specify what they value most within the system message, steering the LLM’s generation behavior to better align with the user’s intentions. However, a naive application of such an approach is non-trivial since LLMs are typically trained on a uniform system message (e.g., "You are a helpful assistant") which limits their ability to generalize to diverse, unseen system messages. To improve this generalization, we create the Multifaceted Collection, a preference dataset with 192k combinations of values beyond generic helpfulness and harmlessness, spanning 65k user instructions. Using this dataset, we train a 7B LLM called Janus and test it on 921 prompts from 5 benchmarks (AlpacaEval 2.0, FLASK, Koala, MT-Bench, and Self-Instruct) by adding various unseen system messages that reflect user preferences. Janus achieves tie+win rate of 75.2%, 72.4%, and 66.4% against Mistral 7B Instruct v0.2, GPT-3.5 Turbo, and GPT-4, respectively. Unexpectedly, on three benchmarks focused on response helpfulness (AlpacaEval 2.0, MT-Bench, Arena Hard Auto v0.1), Janus also outperforms LLaMA 3 8B Instruct by a +4.0%, +0.1%, +3.0% margin, underscoring that training with a vast array of system messages could also enhance alignment to the general public’s preference as well. Our code, dataset, benchmark, and models are available at https://github.com/kaistAI/Janus.
@@ 1 Introduction
Post-training techniques such as reinforcement learning from human feedback (RLHF) or instruction fine-tuning are effective at enhancing the alignment of large language models (LLMs) with human preferences [4, 5, 48, 6, 30, 52]. These methods involve collecting preference data based on high-level values that many people agree upon, such as helpfulness and harmlessness, and then using this data to train reward models (RMs) and LLMs. However, human preferences cannot simply be categorized into such binary divisions; they are typically diverse and exhibit nuanced variations according to different people and contexts [8, 71, 27, 28, 33, 37, 35, 36, 62]. As a result, a significant amount of data is created without considering the conflicting preferences that individuals may have, which results in annotation disagreement [73, 18, 75]. Consequently, LLMs trained on data derived from ambiguous preferences–under the guise of the general public’s preferences, without specifying whose preferences they are–unintentionally develop undesirable traits, such as verbosity [57].
To address these issues, a new paradigm called personalized RLHF has been proposed. In this paradigm, RMs or LLMs are trained to reflect individual preferences, aiming to model the diverse preferences that people may have [71, 8, 21, 7, 35, 39]. However, prior works have a critical drawback: it requires collecting data that reflects fine-grained preferences each time and then retraining the RMs and LLMs. Considering that the size of LLMs used in production is steadily increasing, this method may not be highly scalable in terms of time, storage cost, and training cost [56, 34].
In this paper, we pose the following question: Can we align to diverse preferences without the need to re-train the LLM for each user’s preferences? Drawing inspiration from the transformation of the ‘pretrain-and-finetune’ paradigm to instruction fine-tuning—which allows for handling unseen tasks without re-training [70, 47, 64]—we explore adapating to the user’s preference without re-training. Specifically, we examine how explicitly stating the user’s preference in the system message can guide the LLM’s behavior to align with the user’s intentions [66]. However, in our early experiments, we observed that open-source LLMs, unlike proprietary LLMs, do not possess this ability. We hypothesize that this is because open-source LLMs are typically trained with a unified system message, such as "You are a helpful assistant, " [76, 8, 78, 20, 9]. Consequently, their ability to incorporate new, unseen system message is limited. To address this, we propose training LLMs on diverse system message, allowing the LLM to adapt to novel preferences during test time.
@@
To this end, we introduce the Multifaceted Collection, a preference dataset containing 192k unique system messages for aligning LLMs to diverse human preferences. The Multifaceted Collection consists of 65k unique instructions, each accompanied by 3 system messages, along with respective responses. The instructions are acquired from five existing datasets (Chatbot Arena Conversations [76], Domain-Specific Preference dataset [8], UltraFeedback-binarized-clean [9], Nectar [78], and OpenHermesPreferences [20]). To distill the proprietary LLM’s ability to adhere to the given system message, we generate the system messages and reference answers with GPT-4-Turbo-0125. Through a set of quantitative analyses, we demonstrate that the Multifaceted Collection encompasses a variety of preferences. Then, we train Mistral 7B v0.2 [22] on the Multifaceted Collection and obtain our resulting LLM, Janus 7B.
To test if Janus could adhere to various system messages for alignment, we sample 921 instructions across 5 benchmarks (AlpacaEval 2.0 [13], FLASK [74], Koala [15], MT-Bench [76], and Self-Instruct [69]) and add the additional system messages. The quality of the system messages and reference answers are verified by humans, and instances identified as low-quality are eliminated. In a pairwise ranking human evaluation setting, Janus achieve a 74.8%, 70.8%, and 57.9% win rate compared to Mistral 7B Instruct v0.2, GPT-3.5-Turbo-0125, and GPT-4-Turbo-0125, respectively. Results in direct assessment with GPT-4 evaluation show similar trends, where Janus achieves an average score of 4.24 out of 5.0, outperforming LLaMA 3 8B Instruct (4.1), Mixtral 8x7B Instruct v0.1 (4.03), GPT-3.5-Turbo-0125 (3.91), and closing the gap with LLaMA 3 Instruct 70B (4.39), GPT-4-0613 (4.10), and GPT-4-Turbo-0125 (4.35). Surprisingly, when assessing the helpfulness of the response on 3 benchmarks (AlpacaEval 2.0, MT-Bench, Arena Hard Auto v0.1), Janus outperforms LLaMA 3 8B Instruct by a +4.0%, +0.1%, and +0.3% margin, implying that training on various system messages efficiently enhances alignment to the general public’s preference as well.
@@ 2 Related work
Aligning to diverse preferences
@@
Providing context to an LLM is as easy as giving a textual prompt as input. When a user desires to instill specific behavior when performing a task into an LLM, such as impersonating a role [49, 54] or personalization [43], a prevalent method is to simply include the context as part of the user instruction. The component of the model input specialized for such purpose, coined as system message [66], is introduced with ChatGPT111platform.openai.com/docs/guides/prompt-engineering [47]. For closed-source models accessible by API services, system messages can be set by developers in the API request (e.g., Mistral222docs.mistral.ai/capabilities/guardrailing/, Claude333docs.anthropic.com/en/docs/system-prompts, Command R444docs.cohere.com/docs/preambles). This feature facilitates analysis of the behaviors of top-performing LLMs by diversifying social roles and personalities in system messages [25, 77]. Open-source models report to have trained with specific system messages in an input template include Orca [44] and LLaMA 2 [64], with the purpose of invoking step-by-step explanation or maintaining consistency of responses, respectively. While these works corroborate that diversifying system messages instead of default system messages improves performance, the content of system messages studied in previous works is limited in number and domain, making it insufficient to observe the entire space of human preferences. Similar to scaling instruction to improve LLM’s capability to solve unseen tasks [41], we scale system messages into multifaceted variants, allowing the model to steer generations to be aligned with user’s preferences for the same instruction.
3 Training on various system messages for alignment
3.1 Role of system messages
Existing alignment datasets often vaguely reflect general preferences like helpfulness and harmlessness. We aim to develop a dataset that captures more fine-grained preferences, such as "code-centric style" or "ensuring code ethics" for a coding solution, across multiple dimensions. In this sense, we conceptualize a preference as a detailed textual description of a quality that a desirable response should possess. Firstly, we clarify two requirements for a model to reflect the diversity of human preferences:
@@ 1. R1:
    Multifacetedness: Multiple facets constitute an individual’s preference, e.g., an individual may prefer a response that provide solutions that are immediately applicable, falls under the intermediate level, are varying, and ensures code ethics at the same time, each element indicating a distinct property of a response (See Figure 2). To represent potential facets in preferences of countless individuals, an unbounded space of distinct facets must be guaranteed.
@@ 2. R2:
    Explicitness: Preferences only latently exist in a pair of chosen and rejected responses [58]. To facilitate learning to discriminate between the nuances in the chosen and rejected responses, the rationale of the decision should be made explicit to the model.
Taking the requirements into account, we develop a novel dataset construction approach (Figure 2). To address R1, we devise a hierarchical preference augmentation strategy: starting from the most $n$ general dimensions, $m$ specific subdimensions are branched, and $l$ preferences are formulated under each subdimension. Assuming $n\ll m\ll l$, this structure ensures that a variety of preference facets are defined. Moreover, combining preferences from different dimensions can effectively represent the complex interplay of preferences. To satisfy R2, such multifaceted preferences are exposed in the model input via system messages that are prepended to the instruction. Verbalizing preferences in the input provides more interpretable signals for the model to map preferences to the output.
Figure 2: Construction process of Multifaceted Collection. For each instruction, preference descriptions are augmented from general to specific, allowing for multiple facets to branch out. We combine preferences from various dimensions into a system message to materialize these preferences as model input. Following the system message and instruction, a proprietary LLM generates a gold response for training.
3.2 Data construction and instruction tuning
@@
We first select 65k instructions from a pool of existing 4 high-quality preference datasets: Chatbot Arena Conversations [76], Domain-Specific Preference dataset [8], UltraFeedback-binarized-clean [9], Nectar [78], and OpenHermesPreferences [20]. These datasets curate or rank responses with variants of helpfulness or harmlessness criteria, signifying that the input instructions can be associated with various preferences. We further deduplicate and filter instructions that specifies a preference in a format similar to a system message. The filtering process is detailed in Appendix A.1.
@@
To develop our hierarchical preference augmentation strategy, we start by bootstrapping from manually created seed preferences. Drawing on studies that outline characteristics of language model responses [44, 21, 36], we initially identify four main dimensions for response preferences: style, background knowledge, informativeness, and harmlessness. We then define a preference set consisting of one preference from each dimension. We expand this by brainstorming 18 subdimensions and describing 107 preferences. For each of the 65k instructions, we use GPT-4-Turbo-0125 [1] to generate new, task-aligned preference sets. In each iteration, a randomly chosen preference set from the seeds serves as a model. This process is repeated three times, creating three variations per instruction, totaling 198,000 preference sets. We demonstrate significant variation in preference sets across instructions in Section B.1.
@@
Preference 797,904
System message and gold response generation.
We convert each preference set into a system message using GPT-4 Turbo, generating three system messages per instruction. We manually create 50 system messages and select three random examples from these to ensure comprehensive coverage of all preferences. After a thorough filtering and integrity check, the total number of system messages produced reaches 192k. To generate training data, we use GPT-4 Turbo to craft gold-standard multifaceted responses for each system message. For preference optimization methods that need pairwise inputs, we choose one system message and its corresponding response as the preferred option and one of the two other responses as the rejected option for each instruction. The final dataset, Multifaceted Collection, comprises 192k instances for supervised fine-tuning and 65k instances for preference optimization.
Instruction fine-tuning
@@ 4 Experimental setup
  • Multifacetedness benchmark: Current LLM benchmarks typically assess responses based on general helpfulness, omitting context-specific nuances. To remedy this, we enhance five existing benchmarks–AlpacaEval 2.0 [13], FLASK [74], Koala [15], MT-Bench [76], and Self-Instruct [69]–to evaluate whether a language model can generate responses tailored to specific preferences when provided with context. For convenience, we refer to this enhanced set of benchmarks as Multifaceted Bench. We develop Multifaceted Bench by adapting existing instructions to produce synthetic system messages, aligned reference answers, and a scoring rubric with a proprietary LLM’s assistance. Human annotators are employed to validate the quality of instances. The final test set, with 2.5% substandard samples filtered out from the initial 945 instances, results in 921 instances. Further details on the data curation and filtering process are available in section C.1,  F.1.
  • Helpfulness benchmarks: We utilize three benchmarks, namely Alpaca Eval 2.0 [13] which focuses on single-turn conversations using 805 test prompts; MT-Bench [76] which centers on multi-turn interactions using 160 prompts across eight domains; Arena Hard Auto v0.1 [38] which utilizes 500 prompts collected from the Chatbot Arena platform.
  • Harmlessness benchmark: We utilize RealToxicityPrompts [14] to assess the model’s harmlessness from three perspectives: overall toxicity, fluency, and diversity. Following previous works [40, 42], we employ a test set of 10,000 entries.
@@
We evaluate Janus suite against varying sizes of popular pre-trained open-source models, instruction-tuned open-source models, and preference-optimized proprietary models. Pre-trained models include Mistral 7B v0.2 [22] and LLaMA 3 {8B, 70B} [2]. Instruction-tuned open-source models include LLaMA 2 Chat 70B [64], Mistral 7B Instruct v0.2 [22], Mixtral 8x7B Instruct v0.1 [23], and LLaMA 3 Instruct {8B, 70B} [2]. Proprietary models include GPT-3.5-Turbo-0125 [47], GPT-4-0613 [1], and GPT-4-Turbo-0125. For helpfulness and harmlessness benchmarks, we additionally retrieve results from the official leaderboard or paper.
@@
Our evaluation approach includes assessment by human and LLMs, i.e., LLM-as-a-Judge [76, 26, 33, 28]. The Multifaceted Bench requires that both a system message and an instruction are provided to the model, which then uses these two inputs to generate a response. Given results on Multifaceted Bench, human evaluators compare models pairwise to determine which response is better according to the associated score rubrics. The LLM-as-a-Judge method utilizes a proprietary model, GPT-4-Turbo-0125, to perform absolute scoring of 1 to 5. On helpfulness results, proprietary LLMs judge model responses in direct competition, with evaluations in AlpacaEval 2.0 and MT-Bench scored from 0 to 10, while Arena Hard Auto v0.1 uses pairwise comparisons to calculate an Elo score. To measure toxicity in RealToxicityPrompts, we employ the Perspective API [^5] https://www.perspectiveapi.com/. Detailed evaluation protocols are in Appendix F.

### v2

@@ Abstract
Although humans inherently have diverse values, current large language model (LLM) alignment methods often assume that aligning LLMs with the general public’s preferences is optimal. A major challenge in adopting a more individualized approach to LLM alignment is its lack of scalability, as it involves repeatedly acquiring preference data and training new reward models and LLMs for each individual’s preferences. To address these challenges, we propose a new paradigm where users specify what they value most within the system message, steering the LLM’s generation behavior to better align with the user’s intentions. However, a naive application of such an approach is non-trivial since LLMs are typically trained on a uniform system message (e.g., “You are a helpful assistant”), which limits their ability to generalize to diverse, unseen system messages. To improve this generalization, we create Multifaceted Collection, augmenting 66k user instructions into 197k system messages through hierarchical user value combinations. Using this dataset, we train a 7B LLM called Janus and test it on 921 prompts from 5 benchmarks (AlpacaEval 2.0, FLASK, Koala, MT-Bench, and Self-Instruct) by adding system messages that reflect unseen user values. Janus achieves tie+win rate of 75.2%, 72.4%, and 66.4% against Mistral 7B Instruct v0.2, GPT-3.5 Turbo, and GPT-4, respectively. Unexpectedly, on three benchmarks focused on response helpfulness (AlpacaEval 2.0, MT-Bench, Arena Hard Auto v0.1), Janus also outperforms LLaMA 3 8B Instruct by a +4.0%p, +0.1%p, +3.0%p margin, underscoring that training with a vast array of system messages could also enhance alignment to the general public’s preference as well. Our code, dataset, benchmark, and models are available at https://lklab.kaist.ac.kr/Janus/.
@@ 1 Introduction
Post-training techniques such as reinforcement learning from human feedback (RLHF) or instruction fine-tuning are effective at enhancing the alignment of large language models (LLMs) with human preferences [4, 5, 56, 6, 33, 61]. These methods involve collecting preference data based on high-level values that many people agree upon, such as helpfulness and harmlessness, and then using this data to train reward models (RMs) and LLMs. However, human preferences cannot simply be categorized into such binary divisions; they are typically diverse and exhibit nuanced variations according to different people and contexts [8, 84, 30, 31, 36, 40, 38, 39, 74]. As a result, a significant amount of data is created without considering the conflicting preferences that individuals may have, which results in annotation disagreement [87, 21, 90]. Training LLMs on generic, unattributed preference data inadvertently leads to undesirable traits like verbosity [69].
To address these issues, a new paradigm called personalized RLHF has been proposed. In this paradigm, RMs or LLMs are trained to reflect individual preferences, aiming to model the diverse preferences that people may have [84, 8, 25, 7, 38, 42]. However, prior works require collecting data that reflects fine-grained preferences each time and then re-training the RMs and LLMs. Considering that the size of LLMs used in production is steadily increasing, this method may not be highly scalable in terms of time, storage cost, and training cost [67, 37].
In this paper, we pose the following question: Can we align to diverse preferences without the need to re-train the LLM for each user’s preferences? Drawing inspiration from how pretrain-then-finetune evolved to handle unseen tasks without re-training via instruction fine-tuning [83, 55, 77], we explore to adapt to user preference without re-training. Specifically, we examine how explicitly stating the user’s preference as meta-instructions in the form of system messages can guide the LLM’s behavior to align with the user’s intentions [79]. However, in our early experiments, we observed that open-source LLMs, unlike proprietary LLMs, do not possess this ability. We hypothesize that this is because open-source LLMs are typically trained with a generic system message, such as “You are a helpful assistant” [92, 8, 94, 23, 9]. Consequently, their ability to incorporate new, unseen system message is limited. To address this, we propose training LLMs on diverse system messages, allowing the LLM to adapt to novel preferences during test time.
@@
To acquire sufficiently diverse and multifaceted values that control individual preferences, we devise a hierarchical synthesis of user values from high-level dimensions. Verbalizing combinations of user values into system messages, we create a training dataset of 197k system messages, Multifaceted Collection. For each instruction, there are three variants of system messages embedded with non-overlapping values, along with respective responses. Through quantitative and qualitative analyses, we show that the dataset encompasses a variety of preferences.
To test if our strategy helps aligning to various system messages, we train Mistral 7B v0.2 [26] on the Multifaceted Collection and obtain Janus 7B. In addition, we augment 315 instructions across 5 LLM benchmarks (AlpacaEval 2.0 [13], FLASK [88], Koala [17], MT-Bench [92], and Self-Instruct [82]) with three unseen context-specific system messages verified by humans per instruction, creating 921 unseen system messages. Human evaluation demonstrates that Janus achieve a 74.8%, 70.8%, and 57.9% win rate compared to Mistral 7B Instruct v0.2, GPT-3.5-Turbo-0125, and GPT-4-Turbo-0125, respectively. GPT-4 evaluation results show similar trends, where Janus achieves an average score of 4.24 out of 5.0, outperforming LLaMA 3 8B Instruct, Mixtral 8x7B Instruct v0.1, and GPT-3.5-Turbo-0125 by 5.7% on average. Surprisingly, when assessing the helpfulness of the response on 3 benchmarks (AlpacaEval 2.0 [13], MT-Bench [92], Arena Hard Auto v0.1 [41]), Janus outperforms LLaMA 3 8B Instruct by a +4.0%p, +0.1%p, and +0.3%p margin, suggesting that training on diverse system messages enhances both individualized and general alignment. Further analyses reveal that explicit exposure to diverse preferences during training is crucial, ensuring robust performance even without system messages at test time.
@@ 2 Related work
Personalized RLHF for aligning to diverse preferences
@@
Providing context to an LLM is as easy as giving a textual prompt as input. When a user desires to instill specific behavior when performing a task into an LLM, such as impersonating a role [57, 64] or personalization [49], a prevalent method is to simply include the context as part of the user instruction. The component of the model input specialized for such purpose, coined as system message [79], is introduced with ChatGPT111platform.openai.com/docs/guides/prompt-engineering [55]. For closed-source models accessible by API services, system messages can be set by developers in the API request (e.g., Mistral222docs.mistral.ai/capabilities/guardrailing/, Claude333docs.anthropic.com/en/docs/build-with-claude/prompt-engineering/system-prompts, Command R444docs.cohere.com/docs/preambles). This feature facilitates analysis of the behaviors of top-performing LLMs by diversifying social roles and personalities in system messages [29, 93]. Open-source models report to have trained with specific system messages in an input template include Orca [51] and LLaMA 2 [77], with the purpose of invoking step-by-step explanation or maintaining consistency of responses, respectively. While these works corroborate that diversifying system messages instead of default system messages improves performance, the content of system messages studied in previous works is limited in number and domain, making it insufficient to observe the entire space of human preferences. Similar to scaling instruction to improve LLM’s capability to solve unseen tasks [47], we scale system messages into multifaceted variants, allowing the model to steer generations to be aligned with user’s preferences for the same instruction.
Instruction tuning
Traditional instruction tuning approaches [47, 83, 82, 20, 65, 46] focus on improving model performance across diverse tasks by training on varied instructions. Our work extends this concept by leveraging system messages to encode user preferences, providing a novel approach to individualized alignment. While instruction tuning typically embeds task-specific instructions within user prompts, we utilize system messages to separate preference specifications from task instructions. Concretely, we view system messages as meta-instructions that guide a model how to respond to subsequent instructions, which allows for more nuanced control over model behavior in specific training settings [79]. Our approach curates meta-instructions to specifically simulate user values instead of presenting irrelevant challenges or hinting solutions with respect to the instruction. To the best of our knowledge, there is no work that shares a similar motivation with us or has publicly released a training dataset containing meta-instructions.
3 Multifaceted Collection for scalable individualized alignment
3.1 Mapping a preference to multidimensional values
Existing alignment datasets often vaguely reflect general preferences such as helpfulness and harmlessness. Our aim is to develop a dataset that captures more fine-grained preferences. Specifically, we posit that even for the same instruction, the choice of which response to select or reject (i.e., preference) may differ among individuals due to their unique set of values. For example, given a coding problem, one response focused on code and another centered on explaining concepts might each be chosen as preferable over the other, depending on what the user values as informative. We identify two requirements for a model to effectively reflect such diversity of human preferences:
@@ 1. R1:
    Multifacetedness: Multiple facets can dictate an individual’s preference, not just one. A sufficiently large space of potential facets should be captured in order to model preferences of countless individuals.
@@ 2. R2:
    Explicitness: Preferences only latently exist in a pair of chosen and rejected responses [70]. To facilitate learning to discriminate the nuances in between a chosen response and a rejected response, the rationale of the decision should be made explicit to the model.
Taking the requirements into account, we develop a novel dataset construction approach as illustrated in Figure 2. To address R1, we devise a hierarchical value augmentation strategy: starting from the $n$ most general dimensions, $m$ specific subdimensions are branched, and $l$ values are created under each subdimension. Assuming $n\ll m\ll l$, this structure ensures that a variety of facets are defined. Ultimately combining values from different dimensions, which we define as a preference set, can effectively represent the complex interplay of values to determine a preference. To satisfy R2, we conceptualize a value as a detailed textual description of a quality that a desirable response should possess. Verbalization helps the model become more aware of preference patterns [50] and provides interpretable signals to induce better decisions [68]. The verbalized preference set is contextualized in the model input along with the original instruction given by the user. Precisely, it serves as a meta-instruction which sets a specific direction for the instruction to be executed and hence the additional context is included in the system message prepended to the instruction.
Figure 2: Multifaceted Collection construction process. For each instruction, value descriptions are augmented from general to specific, allowing for multiple facets to branch out. We combine value from various dimensions into a system message to materialize preferences into model input. Following the system message and instruction, a proprietary LLM generates a gold response for training.
3.2 Instruction data construction
Instruction fine-tuning can achieve strong generalization on unseen tasks given a sufficiently large training set of tasks and instructions [65, 46, 30]. To provide a ground for unseen system message generalization and therefore scalable individualized alignment, we implement a LLM-driven data synthesis pipeline to scale the number of multifaceted system messages.
@@ 3.2 Instruction data construction
We first select 66k instructions from a pool of existing four high-quality preference datasets: Chatbot Arena Conversations [92], Domain-Specific Preference dataset [8], UltraFeedback-binarized-clean [9], Nectar [94], and OpenHermesPreferences [23]. These datasets curate or rank responses with variants of helpfulness or harmlessness criteria, signifying that the input instructions can be associated with various preferences. We further deduplicate and filter instructions that specifies a preference in a format similar to a system message.
@@
To implement our hierarchical value augmentation strategy, we start by bootstrapping from manually created seed value descriptions. Drawing on studies that outline characteristics of language model responses [51, 25, 39], we initially identify four main dimensions for response preferences: style, background knowledge, informativeness, and harmlessness. We then brainstorm 18 subdimensions and describe 107 values, additionally taking inspiration from works that delineate user intention in instructions [59], machine personality [28], moral judgment [54], cultural alignment [3], and fine-grained evaluation criteria of language models [30]. For each of the 66k instructions, we use GPT-4-Turbo-0125 [1] to generate a novel, task-aligned preference set given a randomly chosen preference set from the seeds for few-shot learning. This process is repeated three times, creating three varying preference sets per instruction.
@@
Value 797,904
System message and gold response generation
We convert each preference set into a system message using GPT-4-Turbo-0125, generating three system messages per instruction. The model is provided with 3 random examples from our manual creation of 50 system messages to match the format. To generate training data, we use the same LLM to craft gold-standard multifaceted responses for each system message. To create pairwise inputs for preference tuning or reward modeling, we choose one system message and its corresponding response as the preferred option and one of the two other responses as the rejected option for each instruction. The final dataset, Multifaceted Collection, comprises 197k individual responses and 66k pairs of chosen and rejected responses.
We observe low similarity between preference sets for each instruction and demonstrate high quality and safety of system messages in Appendix E. Further details of the construction process are in Appendix D. All prompt templates used for data generation are in Appendix M.
4 Experimental setup
Training
@@ 4 Experimental setup
  • Multifacetedness benchmark: Current LLM benchmarks typically assess responses based on general helpfulness. To evaluate whether a model can generate responses tailored to specific user preferences provided as context, we enhance five benchmarks: AlpacaEval 2.0 [13], FLASK [88], Koala [17], MT-Bench [92], and Self-Instruct [82]. For convenience, we refer to the enhanced set of benchmarks as Multifaceted Bench. We synthesize system messages adapted to existing instructions, corresponding reference answers, and instance-specific scoring rubrics, assisted by a proprietary LLM. Human annotators are employed to validate the quality of instances. The final test set, with 2.5% substandard samples filtered out from the initial 945 instances, results in 921 instances. Further details on the data curation and filtering process are in Appendix G. Train-test similarity and difficulty of instances are reported in Appendix H.
  • Helpfulness benchmarks: We utilize 3 benchmarks, namely 805 test prompts from AlpacaEval 2.0 [13], which focuses on single-turn conversations; 160 prompts across 8 domains in MT-Bench [92], which is centered on multi-turn interactions; 500 prompts from Arena Hard Auto v0.1 [41], which are collected from the Chatbot Arena platform.
  • Harmlessness benchmark: We mainly utilize RealToxicityPrompts [16] to assess harmlessness from three dimensions: overall toxicity, fluency, and diversity. Following previous works [44, 48], we employ a test set of 10k entries. In addition, we evaluate on social bias benchmarks following Team et al. [76]: Winogender [63], CrowS-Pairs [53], and BBQ [58].
@@
We evaluate Janus suite against varying sizes of popular pre-trained open-source models, instruction-tuned open-source models, and preference-optimized proprietary models. Pre-trained models include Mistral 7B v0.2 [26] and LLaMA 3 {8B, 70B} [2]. Instruction-tuned open-source models include LLaMA 2 Chat 70B [77], Mistral 7B Instruct v0.2 [26], Mixtral 8x7B Instruct v0.1 [27], and LLaMA 3 Instruct {8B, 70B} [2]. Proprietary models include GPT-3.5-Turbo-0125 [55], GPT-4-0613 [1], and GPT-4-Turbo-0125. For helpfulness and harmlessness benchmarks, we additionally retrieve results from official leaderboards or previous papers.
@@
We include assessment by both human and LLMs, i.e., LLM-as-a-Judge [92, 30, 36, 31]. Multifaceted Bench requires that both a system message and an instruction are provided to the model to generate a response. Given model responses to Multifaceted Bench, human evaluators compare responses pairwise to determine which response is better, according to the associated score rubrics. A proprietary model, GPT-4-Turbo-0125, also performs absolute scoring of 1 to 5. On helpfulness benchmarks, proprietary LLMs either judge model responses using absolute scoring in the range of 0 to 10 (AlpacaEval 2.0 and MT-Bench) or compare responses pairs to calculate an Elo score (Arena Hard Auto v0.1). To measure toxicity in RealToxicityPrompts, we employ the Perspective API555perspectiveapi.com/. Detailed evaluation protocols are in Appendix K-L.